OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!! A pair of equations is shown below: y = 6x − 4 y = 5x − 3 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations. (6 points) Part B: What is the solution to the pair of equations? (4 points)

OpenStudy (anonymous):

Part A: Lets roughly draw these two line equations without beign precise:  Their intersection will give us the solution to this system (I wasn't precise about drawing). Angle between the first equation and the x-axis will be more than the second's, since first one's slope is greater then the secon one's. Part B: We can write the second equation like this. 6x - 4 = 5x - 3 (We substitute the y in second equation) If we solve for this for x, we would get: x = 1 We put x in one of the equations to find y, we would get: y = 6(1) - 4 = 2